OpenStudy (anonymous):

I*(-1)^11 ...What does this simplify to? and what are the powers of "i"? Also, Can you give me two more similar problems to i^23 and explain how to do it, to get an idea.

OpenStudy (anonymous):

@BangkokGarrett Can you help me?

OpenStudy (anonymous):

Is that an i times -1 raised to the 11th power ???

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

well..(-1)^11 = -1.......-1 ^ any odd number will equal -1 then, i times -1 will just be -i. So the answer is -i .

OpenStudy (anonymous):

i = i i^2 = -1 i^3 = -i i^4 = 1 Then it repeats the same cycle... i^5 = i i^6 = -1 etc etc.

OpenStudy (anonymous):

To find a big one like i^23...find the nearest multiple of 4 24 is a multiple of 4, right? So, we know i^24 = 1....since it goes in cycles of 4 like the pattern I just showed you. Now just count backwards in the cycle to get i^23 We just need to count back 1 .. so i^23 = -i

OpenStudy (anonymous):

oh okaay thankyou!!!

OpenStudy (anonymous):

i^102 100 is the nearest multiple of 4 so i^100 = 1 now just count forward i^101 = i i^102 = -1

OpenStudy (anonymous):

can you give me one more like that please and How do I simplify \[x=-2\pm \sqrt{-6}\]?

OpenStudy (unklerhaukus):

\[\sqrt{-6}=\sqrt{-1\times 6}=\sqrt{-1}\sqrt6=6i\]

OpenStudy (anonymous):

thankyou can you do me one more favor?